### v1

# Vizing’s Conjecture: A Density-Based Re-framing

Abstract
We recast Vizing’s conjecture $\gamma(G\square H)\geq\gamma(G)\gamma(H)$ through a simple domination–density lens. Defining $\rho_{X}=\gamma(X)/|V(X)|$, the conjecture becomes $\rho_{G\square H}\geq\rho_{G}\rho_{H}$. Many classical estimates admit this normalised form, but if we select valid upper bounds $\rho_{G}\leq\tilde{\rho}_{G}$, $\rho_{H}\leq\tilde{\rho}_{H}$ and a product lower bound $\rho_{G\square H}\geq\tilde{\rho}_{G\square H}$, then the single test $\tilde{\rho}_{G\square H}\geq\tilde{\rho}_{G}\tilde{\rho}_{H}$ certifies Vizing’s. Instantiations give: (i) a bipartite imbalance criterion, where partition asymmetry together with a maximum–degree bound yields infinite nontrivial families; and (ii) a regular high–degree regime via the Arnautov–Payan bound, certifying all $k$-regular pairs for $k\geq 32$. The framework is modular: sharper family–specific bounds can be dropped in without altering the proof, expanding the certified regime while preserving the simple density form.
1 Domination Density Framework
Define domination densities:
$\rho_{G}:=\frac{\gamma(G)}{|V(G)|},\quad\rho_{H}:=\frac{\gamma(H)}{|V(H)|},\quad\rho_{G\square H}:=\frac{\gamma(G\square H)}{|V(G)||V(H)|}.$ (1)
Then Vizing’s conjecture is equivalent to:
$\rho_{G\square H}\geq\rho_{G}\rho_{H}.$ (2)
This condition allows us to reason about domination structure irrespective of graph order, providing a continuous framework for the conjecture.
Throughout, let $\Delta(X)$ and $\delta(X)$ denote the maximum and minimum degrees of a graph $X$.
2 Introduction
Vizing’s conjecture (1968) asserts that for all finite simple graphs $G,H$,
$\gamma(G\square H)\ \geq\ \gamma(G)\gamma(H).$
Here $\gamma(\cdot)$ denotes the domination number and $G\square H$ is the Cartesian product. The conjecture has withstood more than five decades of study: it is known to hold for various special classes (trees, cographs, interval graphs, among others), yet no general proof or counterexample has been found. Because domination numbers are notoriously difficult to compute, much of the progress has come from developing upper and lower bounds that can be combined to certify the inequality in restricted regimes.
The density framework introduced above provides a compact way to organise such bounds. This re-framing turns the search for sufficient conditions into a simple inequality between densities, highlighting the structural parameters that drive the conjecture.
In this short note we demonstrate the utility of this perspective through two case studies. First, we derive a bipartite imbalance criterion showing that sufficiently uneven bipartitions guarantee Vizing’s inequality, even for infinite nontrivial families. Second, we show how the framework recovers a regular high–degree regime, where the Arnautov–Payan bound certifies the conjecture for all kk–regular pairs once $k\geq 32$. These examples illustrate how the density lens not only unifies existing arguments but also provides a flexible template into which sharper bounds may be inserted to enlarge the certified regime.
3 Hybrid Sufficient Conditions and Practical Applications
In practical settings regarding Vizing’s conjecture, particularly for algorithmic searches or large-scale analysis, it is often advantageous to work with collections of known domination bounds tailored to specific graph families. That is, under such conditions, one can define sets of computationally light upper and lower domination density bounds, denoted $\mathrm{U}_{G}=\{\rho_{G1},\rho_{G2},\rho_{G3},\dots,\rho_{Gn}\}$ and $\mathrm{L}_{G}=\{\rho_{G1},\rho_{G2},\rho_{G3},\dots,\rho_{Gn}\}$, where each $\rho_{Gi}$ is a valid bound applicable to the structure of $G$.
By evaluating each bound in $\mathrm{U}_{G}$ for a given graph, one can select the tightest available estimate (e.g., the minimal upper bound), and use this to test whether sufficient conditions for Vizing’s Conjecture are satisfied. This hybrid approach—combining theoretical bounds with computational heuristics—may yield practical methods for certifying Vizing’s conjecture across broad graph classes without requiring explicit computation. This framework may also yield theoretical results restricting the parameter space in which a counterexample to the conjecture could possibly exist.
The framework defined explicitly is as follows: Let $G$ and $H$ be simple finite graphs, let $\mathrm{U}_{G}$ and $\mathrm{U}_{H}$ be the sets of upper density bounds with $\tilde{\rho}_{G}$ and $\tilde{\rho}_{H}$ as the minimum element of said sets so that $\rho_{G}\leq\tilde{\rho}_{G}$, and $\rho_{H}\leq\tilde{\rho}_{H}$:
$\tilde{\rho}_{G}:=\min\left\{\mathrm{U_{G}}\right\},\qquad\tilde{\rho}_{H}:=\min\left\{\mathrm{U_{H}}\right\},$ (3)
Similarly let $\mathrm{L_{G\square H}}$ be the set of lower density bounds applicable to the Cartesian product graph of $G$ and $H$ with $\tilde{\rho}_{G\square H}$ as the maximum element so that $\rho_{G\square H}\geq\tilde{\rho}_{G\square H}$:
$\tilde{\rho}_{G\square H}:=\max\left\{\mathrm{L_{G\square H}}\right\},$ (4)
Combining said bounds the unified sufficient condition to prove Vizing’s conjecture becomes:
Proposition 1 (Bound-certified sufficient condition for Vizing).
Let $G$ and $H$ be finite connected graphs. Suppose $\tilde{\rho}_{G},\tilde{\rho}_{H}$ are valid upper bounds on the domination densities of $G$ and $H$, respectively, and $\tilde{\rho}_{G\square H}$ is a valid lower bound on the domination density of $G\square H$. If
$\tilde{\rho}_{G\square H}\ \geq\ \tilde{\rho}_{G}\,\tilde{\rho}_{H}$ (5)
then Vizing’s inequality holds for the pair $(G,H)$; i.e.,
$\gamma(G\square H)\ \geq\ \gamma(G)\,\gamma(H).$
Proof.
By definition of the bounds,
$\gamma(G)\ \leq\ \tilde{\rho}_{G}\,|V(G)|,\qquad\gamma(H)\ \leq\ \tilde{\rho}_{H}\,|V(H)|,\qquad\gamma(G\square H)\ \geq\ \tilde{\rho}_{G\square H}\,|V(G\square H)|.$
Since $|V(G\square H)|=|V(G)|\,|V(H)|$, we obtain
$\gamma(G\square H)\ \geq\ \tilde{\rho}_{G\square H}\,|V(G)|\,|V(H)|\ \geq\ \tilde{\rho}_{G}\,\tilde{\rho}_{H}\,|V(G)|\,|V(H)|\ \geq\ \gamma(G)\,\gamma(H).$
Extensions to this framework arise by incorporating bounds to curated sets associated with specific graph families. With development, this framework could become a powerful computational and theoretical tool for estimating domination numbers. Through the curation of diverse tight bounds, the theoretical parameter regime for potential contradictory graphs tightens, reducing the search space required to confirm Vizing’s conjecture. Such restrictions may aid in future theoretical proofs which could show transformations that expand the safe regime where Vizing’s conjecture is guaranteed.
4 Bipartite graphs: an imbalance sufficient condition
In this section we apply the framework above using simple domination bounds for bipartite graph pairs to define a small but non-trivial parameter regime for when bipartite graphs hold under Vizing’s conjecture.
Let $G$ and $H$ be connected bipartite graphs. Write the bipartition of $G$ as $V(G)=A_{G}\cup B_{G}$ with $A_{G}\cap B_{G}=\varnothing$, and set
$n_{1}(G):=|A_{G}|,\qquad n_{2}(G):=|B_{G}|,\qquad n_{1}(G)\leq n_{2}(G).$
Similarly, $V(H)=A_{H}\cup B_{H}$ with $n_{1}(H):=|A_{H}|$, $n_{2}(H):=|B_{H}|$, and $n_{1}(H)\leq n_{2}(H)$. Thus
$|V(G)|=n_{1}(G)+n_{2}(G),\qquad|V(H)|=n_{1}(H)+n_{2}(H),$
and
$|V(G\square H)|=|V(G)|\,|V(H)|=(n_{1}(G)+n_{2}(G))(n_{1}(H)+n_{2}(H)).$
For connected bipartite $X$, the smaller part dominates $X$, hence
$\gamma(X)\leq\min\{n_{1}(X),n_{2}(X)\}=n_{1}(X).$
Therefore the upper density bounds
$\tilde{\rho}_{G}:=\frac{n_{1}(G)}{n_{1}(G)+n_{2}(G)},\qquad\tilde{\rho}_{H}:=\frac{n_{1}(H)}{n_{1}(H)+n_{2}(H)}$
are valid for all connected bipartite $G,H$.
For any connected graph $X$ we also use the standard bound
$\gamma(X)\ \geq\ \frac{|V(X)|}{\Delta(X)+1},$
so for the Cartesian product we have $\tilde{\rho}_{G\square H}:=1/(\Delta(G)+\Delta(H)+1)$. Applying these bounds, the condition $\tilde{\rho}_{G\square H}\geq\tilde{\rho}_{G}\tilde{\rho}_{H}$ yields:
Lemma 1 (Imbalance–degree criterion).
Let $G,H$ be connected bipartite graphs with bipartition sizes $n_{1}(G)\leq n_{2}(G)$ and $n_{1}(H)\leq n_{2}(H)$. If
$\Bigl{(}1+\frac{n_{2}(G)}{n_{1}(G)}\Bigr{)}\Bigl{(}1+\frac{n_{2}(H)}{n_{1}(H)}\Bigr{)}\;\geq\;\Delta(G)+\Delta(H)+1,$
equivalently
$\frac{n_{2}(G)}{n_{1}(G)}+\frac{n_{2}(H)}{n_{1}(H)}+\frac{n_{2}(G)}{n_{1}(G)}\frac{n_{2}(H)}{n_{1}(H)}\geq\Delta(G)+\Delta(H),$
then Vizing’s inequality holds for the pair $(G,H)$:
$\gamma(G\square H)\;\geq\;\gamma(G)\,\gamma(H).$
Proof.
From the bounds,
$\frac{1}{\Delta(G)+\Delta(H)+1}\ \geq\ \frac{n_{1}(G)}{n_{1}(G)+n_{2}(G)}\cdot\frac{n_{1}(H)}{n_{1}(H)+n_{2}(H)}.$
Multiplying through and rearranging gives exactly the displayed condition, which by the framework proposition implies Vizing’s inequality. ∎
Remark 1 (One–sided threshold).
Fix $\Delta(G)+\Delta(H)$ and $\frac{n_{2}(H)}{n_{1}(H)}$. The condition in Lemma 1 holds whenever
$\frac{n_{2}(G)}{n_{1}(G)}\;\geq\;\frac{\Delta(G)+\Delta(H)+1}{1+\frac{n_{2}(H)}{n_{1}(H)}}-1.$
In particular, $\frac{n_{2}(G)}{n_{1}(G)}\geq\Delta(G)+\Delta(H)$ or $\frac{n_{2}(H)}{n_{1}(H)}\geq\Delta(G)+\Delta(H)$ suffices, yielding infinite families via the constructions listed below.
Given the limitations of the parameter regime, we list constructions that yield arbitrarily large, structurally rich bipartite graphs which satisfy the sufficient condition. Note that this is easiest to satisfy when $G$ and $H$ are very large with sufficiently uneven bipartitions, i.e. when $n_{1}\ll n_{2}$. Constructions include:
  1. 1.
    One-sided blow ups via leaves and path tree-like structures. Given any initial bipartite core, the addition of one-sided leaves, longer pendant paths, as well as other structures which increase bipartition unevenness push a graph toward the successful parameter regime. Such operations can be repeated indefinitely such that the bipartition becomes sufficiently uneven, i.e. $n_{1}\ll n_{2}$.
  2. 2.
    Even-cycle grafting. Begin with a bipartite graph with an uneven partition. Select a vertex from the larger side and graft an even cycle $C_{2m}$ by identifying one of its vertices with the host. The cycle contributes evenly to both sides, so attaching it to the larger side preserves the imbalance while adding structural complexity (extra cycles, greater diameter, non-tree-like form). Repeating at multiple sites yields dense bipartite families rich in even cycles, yet still uneven, ensuring $\tilde{\rho}_{G\square H}\geq\tilde{\rho}_{G}\tilde{\rho}_{H}$ holds.
5 Regular graphs: a high–degree regime
Applying the same framework to regular graphs we first note that for a $k$-regular graph $X$ we have $\delta(X)=\Delta(X)=k$. We also recall that for Cartesian products,
$\Delta(G\square H)=\Delta(G)+\Delta(H).$
Theorem 1 (Regular-degree threshold).
Let $G$ and $H$ be connected $k$-regular graphs. If $k\geq 32$, then Vizing’s conjecture holds for the pair $(G,H)$, i.e.
$\gamma(G\square H)\ \geq\ \gamma(G)\,\gamma(H).$
Proof.
For $k$-regular $G,H$ we have $\Delta(G\square H)=\Delta(G)+\Delta(H)=2k$, hence
$\gamma(G\square H)\ \geq\ \frac{|V(G)||V(H)|}{2k+1}.$
By the Arnautov–Payan bound (for $\delta\geq 1$),
$\gamma(X)\ \leq\ \frac{H_{\delta(X)+1}}{\delta(X)+1}\,|V(X)|,\qquad H_{m}=\sum_{j=1}^{m}\frac{1}{j}.$
Thus for $k$-regular $G,H$ it suffices that
$\frac{1}{2k+1}\ \geq\ \Bigl{(}\frac{H_{k+1}}{k+1}\Bigr{)}^{\!2}.$
A direct check at $k=32$ gives $1/(65)\approx 0.015385$ while $(H_{33}/33)^{2}\approx 0.015352$, so the inequality holds there. For $k\geq 33$, set n=k+1(34)n=k+1\ ($\geq$ 34) and use the standard estimate $H_{n}\leq\ln n+\gamma+\frac{1}{2n}$. Then
$\Bigl{(}\frac{H_{k+1}}{k+1}\Bigr{)}^{\!2}\ \leq\ \Bigl{(}\frac{\ln n+\gamma+\tfrac{1}{2n}}{n}\Bigr{)}^{\!2}\ \leq\ \Bigl{(}\frac{\ln 34+\gamma+\tfrac{1}{68}}{n}\Bigr{)}^{\!2}\ <\ \frac{1}{2n}\ \leq\ \frac{1}{2k+1},$
since $\ln 34\approx 3.526$, $\gamma\approx 0.5772$, and $2(\ln 34+\gamma+\tfrac{1}{68})^{2}<34$. Hence the inequality holds for all $k\geq 32$. ∎
Remark.
Any refinement of either (i) the product-side bound $\gamma(G\square H)\geq|V(G)||V(H)|/(\Delta(G)+\Delta(H)+1)$ for (near-)regular inputs, or (ii) the Arnautov–Payan upper bound, immediately lowers the $k$-threshold. The utility is plug-and-play: better constants or structure-sensitive estimates (spectral/expander, girth, bipartite) transfer mechanically to sharper high–degree regimes and extend verbatim to bi-regular and almost-regular pairs.
6 Conclusion
We have presented a domination–density framework that re-frames Vizing’s conjecture as a simple inequality between normalised domination numbers. Applying this lens certifies two broad regimes: (i) a bipartite imbalance criterion guaranteeing the conjecture for infinitely many uneven families, and (ii) a high–degree regime where the Arnautov–Payan bound establishes the result for all kk–regular pairs with $k\geq 32$.
The approach is deliberately modular: any tighter upper or lower bounds can be inserted without altering the proof pattern, enlarging the certified space while preserving the clarity of the density form. Future directions include refining the regular threshold via sharper estimates, developing mixed-degree criteria for irregular graphs, and creating algorithmic methods to automatically certify large instances. Overall, the density perspective offers a unifying framework and a path toward narrowing the search for counterexamples.