### v2

# Vizing’s Conjecture: A Density-Based Re-framing

Abstract
We present an equivalent form of Vizing’s conjecture $\gamma(G\square H)\geq\gamma(G)\gamma(H)$ using a simple domination–density lens. Defining $\rho_{G}=\gamma(G)/|V(G)|$, the conjecture becomes $\rho_{G\square H}\geq\rho_{G}\rho_{H}$, that is the dominating density of the Cartesian product is equal to or larger than the product of the domination densities of the original graphs. Selecting valid upper bounds such $\rho_{G}\leq\tilde{\rho}_{G}$, $\rho_{H}\leq\tilde{\rho}_{H}$ and a product lower bound such that $\rho_{G\square H}\geq\tilde{\rho}_{G\square H}$, then we note the simple test $\tilde{\rho}_{G\square H}\geq\tilde{\rho}_{G}\tilde{\rho}_{H}$ which certifies Vizing’s if holds. Examples of this give Vizing’s holds for: (i) bipartite graph pairs whose bi-partitions are sufficiently uneven compared to the graphs maximum degree, yielding infinite nontrivial graph pairs; and (ii) via the Arnautov–Payan bound for all $k$-regular pairs for $k\geq 32$. The framework utilised is modular and the bounds included quite general. The implementation of increasingly sharp family–specific bounds can be utilised to expand the certified parameter regime’s for when Vizing’s holds.
1 Domination Density Framework
Let $\Delta(G)$ and $\delta(G)$ denote the maximum and minimum degrees of a graph $G$. Define domination densities:
$\rho_{G}:=\frac{\gamma(G)}{|V(G)|},\quad\rho_{H}:=\frac{\gamma(H)}{|V(H)|},\quad\rho_{G\square H}:=\frac{\gamma(G\square H)}{|V(G)||V(H)|}.$ (1)
Which implies Vizing’s conjecture is equivalent to:
$\rho_{G\square H}\geq\rho_{G}\rho_{H}.$ (2)
This condition allows us to reason about domination structure irrespective of graph order, providing a continuous framework for the conjecture.
2 Introduction
Vizing’s conjecture (1968) asserts that for all finite simple graphs $G,H$,
$\gamma(G\square H)\ \geq\ \gamma(G)\gamma(H).$
Here $\gamma(G)$ denotes the domination number and $G\square H$ is the Cartesian product. The conjecture is known to hold for various special classes (trees, cographs, interval graphs, among others), yet no general proof or counterexample has been found. Because domination numbers are difficult to compute, much progress has come from developing upper and lower bounds that can be combined to certify the inequality in restricted regimes.
We demonstrate the utility of this via two cases: First, we derive a bipartite regime showing that sufficiently uneven bi-partitions guarantee Vizing’s inequality for infinite nontrivial graph families. Second, we use the Arnautov–Payan bound to certify the conjecture for all kk–regular pairs $k\geq 32$. These illustrate how the density lens not only unifies existing arguments but also provides a flexible template into which sharper bounds may be inserted to enlarge the certified regime.
3 Hybrid Sufficient Conditions and Practical Applications
In practical settings regarding Vizing’s conjecture it is often advantageous to work with collections of known domination bounds tailored to specific graph families. That is, one can define sets of computationally light upper and lower domination density bounds which we denote $\mathrm{U}(G)=\{\rho_{G1},\rho_{G2},\rho_{G3},\dots,\rho_{Gn}\}$ and $\mathrm{L}_{G}=\{\rho_{G1},\rho_{G2},\rho_{G3},\dots,\rho_{Gn}\}$, where each $\rho_{Gi}$ is a valid lower of upper bound applicable to the structure of $G$.
By evaluating element in $\mathrm{U}(G)$ for a given graph, one can select the minimal upper bound and use this to test if sufficient conditions for Vizing’s Conjecture are satisfied. This hybrid approach combining theoretical bounds with computational heuristics may yield theoretical results restricting the parameter space in which a counterexample could possibly exist.
The framework is as follows: Let $G$ and $H$ be simple finite graphs. Let $\mathrm{U}(G)$ and $\mathrm{U}(H)$ be the sets of upper density bounds with $\tilde{\rho}_{G}$ and $\tilde{\rho}_{H}$ as the minimum element so that $\rho_{G}\leq\tilde{\rho}_{G}$, and $\rho_{H}\leq\tilde{\rho}_{H}$:
$\tilde{\rho}_{G}:=\min\left\{\mathrm{U}(G)\right\},\qquad\tilde{\rho}_{H}:=\min\left\{\mathrm{U}(H)\right\},$ (3)
Similarly let $\mathrm{L}(G\square H)$ be the set of lower density bounds applicable to the Cartesian product graph of $G$ and $H$ with $\tilde{\rho}_{G\square H}$ as the maximum element so that $\rho_{G\square H}\geq\tilde{\rho}_{G\square H}$:
$\tilde{\rho}_{G\square H}:=\max\left\{\mathrm{L}(G\square H)\right\},$ (4)
Combining said bounds the unified sufficient condition to prove Vizing’s conjecture becomes:
Proposition 1 (Bound-certified sufficient condition for Vizing).
Let $G$ and $H$ be finite connected graphs. Suppose $\tilde{\rho}_{G},\tilde{\rho}_{H}$ are valid upper bounds on the domination densities of $G$ and $H$, respectively, and $\tilde{\rho}_{G\square H}$ is a valid lower bound on the domination density of $G\square H$. If
$\tilde{\rho}_{G\square H}\ \geq\ \tilde{\rho}_{G}\,\tilde{\rho}_{H}$ (5)
then Vizing’s inequality holds for the pair $(G,H)$; i.e.,
$\gamma(G\square H)\ \geq\ \gamma(G)\,\gamma(H).$
Proof.
By definition of the bounds,
$\gamma(G)\ \leq\ \tilde{\rho}_{G}\,|V(G)|,\qquad\gamma(H)\ \leq\ \tilde{\rho}_{H}\,|V(H)|,\qquad\gamma(G\square H)\ \geq\ \tilde{\rho}_{G\square H}\,|V(G\square H)|.$
Since $|V(G\square H)|=|V(G)|\,|V(H)|$, we obtain
$\gamma(G\square H)\ \geq\ \tilde{\rho}_{G\square H}\,|V(G)|\,|V(H)|\ \geq\ \tilde{\rho}_{G}\,\tilde{\rho}_{H}\,|V(G)|\,|V(H)|\ \geq\ \gamma(G)\,\gamma(H).$
Corollary 1 (Order–invariance of domination density).
Domination densities are invariant under graph order. Consequently, when averaging over the induced sub-graphs of $G\square H$, the mean domination density satisfies
$\overline{\rho}_{\,G\square H}\;\geq\;\rho_{G}\,\rho_{H}.$
Proof.
By definition $\rho_{G}=\gamma(G)/|V(G)|$ and $\rho_{H}=\gamma(H)/|V(H)|$, both of which are ratios independent of scaling in $|V(G)|,|V(H)|$. Thus, densities are order invariant. For each induced sub-graph of $G\square H$, the same density property holds, and by averaging over all such sub-graphs the resulting mean density cannot be less than the product $\rho_{G}\rho_{H}$, which is the lower threshold enforced by Vizing’s inequality. ∎
Through the curation of diverse tight bounds, the theoretical parameter regime for potential contradictory graphs tightens, reducing the search space to confirm Vizing’s conjecture. Such restrictions may aid in future theoretical proofs which could show transformations that expand the safe regime where Vizing’s conjecture is guaranteed.
4 Bipartite Graphs: Bi-partition imbalance regime
In the following section we combine the domination density framework with simple domination bounds applicable to bipartite graphs to derive a bi-partition imbalance regime for when pairs of bipartite graphs satisfy Vizing’s conjecture. Let $G$ and $H$ be connected bipartite graphs. Write the bi-partition of $G$ as $V(G)=A_{G}\cup B_{G}$ with $A_{G}\cap B_{G}=\emptyset$, and set $n_{1}(G):=|A_{G}|$, $n_{2}(G):=|B_{G}|$, and $n_{1}(G)\leq n_{2}(G).$ Similarly for $H$, $V(H)=A_{H}\cup B_{H}$ with $n_{1}(H):=|A_{H}|$, $n_{2}(H):=|B_{H}|$, and $n_{1}(H)\leq n_{2}(H)$. Thus
$|V(G)|=n_{1}(G)+n_{2}(G)\qquad\text{and}\qquad|V(H)|=n_{1}(H)+n_{2}(H)$
where
$|V(G\square H)|=|V(G)|\,|V(H)|=(n_{1}(G)+n_{2}(G))(n_{1}(H)+n_{2}(H)).$
We note that for any connected bipartite graph $G$ the nodes of the smaller bi-partition always dominate $G$, hence
$\gamma(G)\leq n_{1}(G)$
which allows us to define the following bounds which are valid for all connected bipartite graphs
$\rho_{G}\leq\frac{n_{1}(G)}{n_{1}(G)+n_{2}(G)}\qquad\text{and}\qquad\rho_{H}\leq\frac{n_{1}(H)}{n_{1}(H)+n_{2}(H)}.$
Let $\Delta(G)$ and $\delta(G)$ denote the maximum and minimum degrees of a graph $G$. We then note that any connected graph $G$ satisfies the standard bound
$\gamma(G)\ \geq\ \frac{|V(G)|}{\Delta(G)+1},$
so for the Cartesian product, given $\Delta(G\square H)=\Delta(G)+\Delta(H)$, we have
$\rho_{G\square H}\geq\frac{1}{\Delta(G)+\Delta(H)+1}$
Applying these bounds, the condition $\rho_{G\square H}\geq\rho_{G}\rho_{H}$ yields:
Theorem 1 (Imbalance–degree criterion).
Let $G,H$ be connected bipartite graphs with bi-partition sizes $n_{1}(G)\leq n_{2}(G)$ and $n_{1}(H)\leq n_{2}(H)$. If
$\Bigl{(}1+\frac{n_{2}(G)}{n_{1}(G)}\Bigr{)}\Bigl{(}1+\frac{n_{2}(H)}{n_{1}(H)}\Bigr{)}\;\geq\;\Delta(G)+\Delta(H)+1,$
then Vizing’s inequality holds for the pair $(G,H)$, i.e. $\gamma(G\square H)\geq\gamma(G)\gamma(H)$.
Proof.
From the bounds,
$\frac{1}{\Delta(G)+\Delta(H)+1}\ \geq\ \frac{n_{1}(G)}{n_{1}(G)+n_{2}(G)}\cdot\frac{n_{1}(H)}{n_{1}(H)+n_{2}(H)}.$
Multiplying through and rearranging gives exactly the displayed condition, which by the framework proposition implies Vizing’s inequality. ∎
Remark 1 (One–sided threshold).
Fix $\Delta(G)+\Delta(H)$ and $\frac{n_{2}(H)}{n_{1}(H)}$. The condition in Theorem 1 holds whenever
$\frac{n_{2}(G)}{n_{1}(G)}\;\geq\;\frac{\Delta(G)+\Delta(H)+1}{1+\frac{n_{2}(H)}{n_{1}(H)}}-1.$
In particular, $\frac{n_{2}(G)}{n_{1}(G)}\geq\Delta(G)+\Delta(H)$ or $\frac{n_{2}(H)}{n_{1}(H)}\geq\Delta(G)+\Delta(H)$ suffices.
5 Regular graphs: a high–degree regime
Applying the same framework to regular graphs we first note that for a $k$-regular graph $G$ we have $\delta(G)=\Delta(G)=k$. We also recall that for Cartesian products,
$\Delta(G\square H)=\Delta(G)+\Delta(H).$
Theorem 2 (Regular-degree threshold).
Let $G$ and $H$ be connected $k$-regular graphs. If $k\geq 32$, then Vizing’s conjecture holds for the pair $(G,H)$, i.e.
$\gamma(G\square H)\ \geq\ \gamma(G)\,\gamma(H).$
Proof.
For $k$-regular $G,H$ we have $\Delta(G\square H)=\Delta(G)+\Delta(H)=2k$, hence
$\gamma(G\square H)\ \geq\ \frac{|V(G)||V(H)|}{2k+1}.$
By the Arnautov–Payan bound (for $\delta\geq 1$),
$\gamma(G)\ \leq\ \frac{H_{\delta(G)+1}}{\delta(G)+1}\,|V(G)|,\qquad H_{G}=\sum_{j=1}^{G}\frac{1}{j}.$
Thus for $k$-regular $G,H$ it suffices that
$\frac{1}{2k+1}\ \geq\ \Bigl{(}\frac{H_{k+1}}{k+1}\Bigr{)}^{\!2}.$
A direct check at $k=32$ gives $1/(65)\approx 0.015385$ while $(H_{33}/33)^{2}\approx 0.015352$, so the inequality holds there. For $k\geq 33$, set n=k+1(34)n=k+1\ ($\geq$ 34) and use the standard estimate $H_{n}\leq\ln n+\gamma+\frac{1}{2n}$. Then
$\Bigl{(}\frac{H_{k+1}}{k+1}\Bigr{)}^{\!2}\ \leq\ \Bigl{(}\frac{\ln n+\gamma+\tfrac{1}{2n}}{n}\Bigr{)}^{\!2}\ \leq\ \Bigl{(}\frac{\ln 34+\gamma+\tfrac{1}{68}}{n}\Bigr{)}^{\!2}\ <\ \frac{1}{2n}\ \leq\ \frac{1}{2k+1},$
since $\ln 34\approx 3.526$, $\gamma\approx 0.5772$, and $2(\ln 34+\gamma+\tfrac{1}{68})^{2}<34$. Hence the inequality holds for all $k\geq 32$. ∎
Remark.
Any refinement of either (i) the product-side bound $\gamma(G\square H)\geq|V(G)||V(H)|/(\Delta(G)+\Delta(H)+1)$ for (near-)regular inputs, or (ii) the Arnautov–Payan upper bound, immediately lowers the $k$-threshold. The utility is plug-and-play: better constants or structure-sensitive estimates (spectral/expander, girth, bipartite) transfer mechanically to sharper high–degree regimes and extend verbatim to bi-regular and almost-regular pairs.
6 Conclusion
We have presented a domination–density framework that re-frames Vizing’s conjecture as a simple inequality between normalised domination numbers. Applying this lens certifies two broad regimes: (i) a bipartite imbalance criterion guaranteeing the conjecture for infinitely many uneven families, and (ii) a high–degree regime where the Arnautov–Payan bound establishes the result for all kk–regular pairs with $k\geq 32$.
The approach is deliberately modular: any tighter upper or lower bounds can be inserted without altering the proof pattern, enlarging the certified space while preserving the clarity of the density form. Future directions include refining the regular threshold via sharper estimates, developing mixed-degree criteria for irregular graphs, and creating algorithmic methods to automatically certify large instances. Overall, the density perspective offers a unifying framework and a path toward narrowing the search for counterexamples.